# Spin noise spectroscopy of an alignment-based atomic magnetometer

### v1

@@ Abstract
Optically pumped magnetometers (OPMs) are revolutionising the task of magnetic-field sensing due to their extremely high sensitivity combined with technological improvements in miniaturisation which have led to compact and portable devices. OPMs can be based on spin-oriented or spin-aligned atomic ensembles which are spin-polarized through optical pumping with circular or linear polarized light, respectively. Characterisation of OPMs and the dynamical properties of their noise is important for applications in real-time sensing tasks. In our work, we experimentally perform spin noise spectroscopy of an alignment-based magnetometer. Moreover, we propose a stochastic model that predicts the noise power spectra exhibited by the device when, apart from the strong magnetic field responsible for the Larmor precession of the spin, white noise is applied in the perpendicular direction aligned with the pumping-probing beam. By varying the strength of the noise applied as well as the linear-polarisation angle of incoming light, we verify the model to accurately predict the heights of the Larmor-induced spectral peaks and their corresponding line-widths. Our work paves the way for alignment-based magnetometers to become operational in real-time sensing tasks.
@@ I Introduction
In an optical magnetometer, the atoms are spin-polarized using light through the process of optical pumping [26]. In an orientation-based optical magnetometer, the atoms are optically pumped with circularly polarised light. In this case, each atom can be effectively treated as a spin-1/2 particle, even if the ground state of the atom has a total angular momentum $F$ larger than 1/2. The evolution of the atomic spin in a magnetic field is then well described by the Bloch equation for the spin vector $\bm{F}=\left(F_{x},F_{y},F_{z}\right)^{\mathrm{T}}$. Its three components correspond to the expectation values of the angular momentum operators defined along the respective directions, i.e. $F_{\alpha}\coloneqq\mathrm{Tr}\{\hat{F}_{\alpha}\rho\}$ with $\alpha=x,y,z$ for an atomic ensemble being effectively described by a single-atom density matrix $\rho$. On the contrary, in an alignment-based magnetometer [27, 24, 28, 29, 30, 25] the atoms are optically pumped with linearly polarised light. In this case, each atom can be effectively treated as a spin-1 particle [31]. As a result, one has to abandon describing the atomic state with a three-component vector, and instead describe it using rank-2 spherical tensors with five components, which describe how the atomic spin is aligned along certain axes [32].
The purpose of spin noise spectroscopy (SNS) [33] is to characterise the noise properties of a given atomic system and, in particular, the form of the autocorrelation function that noise fluctuations exhibit in the steady-state regime [34]. However, only in the case of orientation-based magnetometers have stochastic noise models been proposed that are capable of explaining the observed noise power spectra when probing unpolarised atomic ensembles [35, 36, 37, 38]. In contrast, such models characterising fully the spin-noise spectra in alignment-based magnetometers are still missing, despite recent promising steps in that direction [39, 40, 41, 42, 43, 44].
@@ II Alignment-based atomic magnetometer
In Fig. 1 we depict the natural spatial configuration of an alignment-based atomic magnetometer [24], which applies to the implementation considered here. A strong magnetic field $\bm{B}_{0}$, directed along $z$, acts perpendicularly to the direction of light propagation, here chosen to be $x$, with light being linearly polarised at an angle $\theta$ to $z$-axis in the $yz$-plane. Both the field and the light interact with an atom and modify its steady state, otherwise induced solely by the relaxation processes [27]. Although we defer the formal description of the atomic steady-state, let us note for now that its form can be conveniently represented by the angular-momentum probability surface [48, 32] (depicted in blue in Fig. 1). As the surface describes the effective polarisation of the steady state, it indicates the rotation symmetries the state possesses. Its shape strongly depends on the polarisation angle $\theta$ of the input light-beam, e.g., resembling a “peanut” for $\theta=0$ or a “doughnut” for $\theta=\tfrac{\pi}{2}$, see Fig. 1(a) and (b), respectively. Importantly, in our experiment the atom is further disturbed by a stochastic field $\bm{B}_{\mathrm{noise}}$, see Fig. 1(c), that induces white noise along the $x$-direction, whose impact can then be monitored by inspecting the polarisation angle $\theta^{\prime}$ of the transmitted beam, while the atomic state is constantly “kicked out” of equilibrium.
Furthermore, as we carry out the experiment in conditions in which correlations between distinct atoms may be neglected during the sensing process, the dynamics of the whole ensemble is effectively captured by the evolution of a single atom [26]. Due to the effective atomic density being low, the shot noise of the photon-detection process dominates over the atomic projection noise, and the impact of the measurement backaction on the atomic steady-state can be ignored [49]. Moreover, thanks to the presence of a relatively strong magnetic field $\bm{B}_{0}$, given that the experiment is carried out at the room temperature, the mechanism of spin-exchange interactions during atomic collisions is irrelevant [50], in contrast to SERF magnetometers [51], that could in principle induce not only classical correlations but also entanglement between individual atoms [52, 53].
@@ II.1 Spherical tensor representation of the atomic steady state
$\rho=\sum_{M,M^{\prime}=-F}^{F}\rho_{MM^{\prime}}^{(F)}\left|F,M\phantom{{}^{\prime}}\right\rangle\!\!\left\langle F,M^{\prime}\right|.$ (1)
However, it is much more convenient to decompose the density matrix in the basis of spherical tensor operators of rank $\kappa=0,1,\dots,2F$, which transform independently in each $\kappa$-subspace under rotations, and hence magnetic fields, in a well-behaved manner. In particular, we may equivalently write any state (1) as
where the density matrix is now decomposed into a sum of rank-$\kappa$ components, each constituting a sum (over $q$) of, in principle, non-Hermitian tensor operators ${{\mathcal{T}}}^{(\kappa)}_{q}$ multiplied by their corresponding complex coefficients $m_{\kappa q}$. Dealing with an atom of fixed $F$, we drop the explicit $F$-dependence of the terms in Eq. (2). As ${{\mathcal{T}}}^{(\kappa)}_{q}$ form a basis, they formally satisfy $\mathrm{Tr}\!\left\{{{\mathcal{T}}}^{(\kappa)}_{q}({{\mathcal{T}}}^{(\kappa^{\prime})}_{q^{\prime}})^{\dagger}\right\}=\delta_{\kappa\kappa^{\prime}}\delta_{qq^{\prime}}$, while it is convenient to also impose ${{\mathcal{T}}}^{(\kappa){\dagger}}_{q}=(-1)^{q}{{\mathcal{T}}}^{(\kappa)}_{-q}$, so that conditions $m_{\kappa q}=(-1)^{q}m^{*}_{\kappa,-q}$ and $\mathrm{Im}\{m_{\kappa 0}\}=0$ assure then the density matrix (2) to be Hermitian. The $m_{00}=1/\sqrt{2F+1}$ coefficient is fixed by the $\mathrm{Tr}\!\left\{\rho\right\}=1$ condition, while the corresponding ($\kappa=0$) tensor operator, $1{{\mathcal{T}}}^{(0)}_{0}=\tfrac{1}{\sqrt{2F+1}}\openone_{2F+1}$, is the only one with a non-zero trace, being invariant under any rotations.
@@
$\rho=\bigoplus_{F}p_{F}\;\rho^{(F)}.$ (3)
Moreover, one may then focus on the dynamics of only one particular $\rho^{(F)}$ for the $F$-level actually contributing to the light-atom interaction, with $\rho^{(F)}$ being then decomposable just as in Eq. (2). In what follows, we drop the $(F)$-superscript for simplicity.
As the case of a linearly polarised pump is of our interest, see Fig. 1, from symmetry arguments it follows that, independently of the spin-number $F$, only even-$\kappa$ coefficients are modified when interacting with light [54]. Moreover, as we show later, our model predicts no significant coupling between components of different $\kappa$. On the other hand, it is only the orientation ($\kappa=1$) and alignment ($\kappa=2$) components that can be probed by resorting to electric dipole light-atom interactions [26, 54]. Hence, even though the pump in the experiment is relatively strong, so that $\rho$ in Eq. (2) (for a particular $F$) reads:
@@
we may assume that all the information, which is accessible when probing the atomic state $\rho$ with light, is in fact completely determined by the vector containing the alignment coefficients:
$\textbf{m}=\left(\begin{array}[]{ccccc}m_{2,-2}&m_{2,-1}&m_{2,0}&m_{2,1}&m_{2,2}\end{array}\right)^{\mathrm{T}}.$ (5)
@@
where $\mathsf{s}_{\theta}\coloneqq\sin\theta$. However, as the strong static field $\bm{B}_{0}$ leads to fast (Larmor) precession of the atomic state around $z$, see Fig. 1, i.e. with the Larmor (angular) frequency $\Omega_{L}\coloneqq\gamma_{\mathrm{gmr}}|\bm{B}_{0}|$ much greater than the overall relaxation rate, all the multipoles $m_{2,q\neq 0}$ quickly average to zero, so that according to the secular approximation [26] the steady-state vector (6) can then be simplified to
@@ II.2.1 Atomic stochastic dynamics
Deterministic noisy evolution.
In general, any noisy evolution of the atomic state (1) is described by the Gorini–Kossakowski–Sudarshan–Lindblad equation [55]:
@@
$B_{\mathrm{noise}}(t)=\frac{\Omega_{\mathrm{noise}}}{\gamma_{\mathrm{gmr}}}\,\xi(t)\;\approx\;\frac{\Omega_{\mathrm{noise}}}{\gamma_{\mathrm{gmr}}}\sqrt{\frac{1}{2\pi f_{\textrm{cutoff}}}}\frac{\mathrm{d}W_{t}}{\mathrm{d}t},$ (15)
where $\xi(t)$ represents the stochastic process for the noise we generate, see App. C for its further characteristics, which effectively exhibits a constant power spectrum that is cut-off at some frequency $f_{\textrm{cutoff}}$ (in Hz). Now, as we will deal with processes occurring at (angular) frequencies $\omega\ll 2\pi f_{\textrm{cutoff}}$, we approximate such noise in Eq. (15) as being white with the correct rescaling factor—the white noise can be interpreted as the time-derivative, $\frac{\mathrm{d}W}{\mathrm{d}t}$, of the Wiener process, $W_{t}$ [56]. As a result, we can write the (stochastic) time-increment induced by the noise involving the Hamiltonian as
@@
where we define $\omega_{\mathrm{noise}}\coloneqq\Omega_{\mathrm{noise}}^{2}/(2\pi f_{\textrm{cutoff}})$ as the effective noise spectral density, while $\mathrm{d}W_{t}\sim\mathcal{N}(0,\mathrm{d}t)$ is then the Wiener increment [56].
@@ II.2.2 Spherical-tensor representation
Formally, this corresponds to the assumption that in the absence of stochastic noise one should differentiate only between dephasing rates along, $\Gamma_{\parallel}\coloneqq\Gamma_{z}$, and perpendicular to, $\Gamma_{\perp}\coloneqq\Gamma_{x}=\Gamma_{y}$, the static field, while keeping the isotropic depolarisation rate, $\Gamma_{\mathrm{iso}}$, as an independent parameter. The effective rates $\bm{\Gamma}$, see also App. A, are then defined by reparametrising the problem as in Eq. (23).
@@
$\mathrm{d}\textbf{m}_{t}$ $=(\textbf{A}_{0}+\frac{\textbf{A}_{\mathrm{noise}}^{2}}{2})\textbf{m}_{t}\,\mathrm{d}t+\textbf{A}_{\Phi}(\textbf{m}_{t}-\textbf{m}^{\mathrm{ss}})\,\mathrm{d}t$
@@ II.2.3 Detected signal
We depict the impact of the Faraday effect in Fig. 1(c), owing to which the angle $\theta$ of the linearly polarised incoming beam is changed to $\theta^{\prime}$ upon leaving the atomic cell. Treating the atomic ensemble as an optically thin sample, the change of the angle, $\Delta\theta\coloneqq\theta^{\prime}-\theta$, obeys then [32]:
@@
where $g_{D}$ is the effective proportionality constant, whereas $\zeta(t)$ denotes the detection stochastic noise, which is completely uncorrelated from the magnetic-field noise affecting the atom dynamics in Eq. (24).

### v2

@@ Abstract
Optically pumped magnetometers (OPMs) are revolutionising the task of magnetic-field sensing due to their extremely high sensitivity combined with technological improvements in miniaturisation which have led to compact and portable devices. OPMs can be based on spin-oriented or spin-aligned atomic ensembles which are spin-polarised through optical pumping with circular or linear polarised light, respectively. Characterisation of OPMs and the dynamical properties of their noise is important for applications in real-time sensing tasks. In our work, we experimentally perform spin noise spectroscopy of an alignment-based magnetometer. Moreover, we propose a stochastic model that predicts the noise power spectra exhibited by the device when, apart from the strong magnetic field responsible for the Larmor precession of the spin, white noise is applied in the perpendicular direction aligned with the pumping-probing beam. By varying the strength of the noise applied as well as the linear-polarisation angle of incoming light, we verify the model to accurately predict the heights of the Larmor-induced spectral peaks and their corresponding linewidths. Our work paves the way for alignment-based magnetometers to become operational in real-time sensing tasks.
@@ I Introduction
In an optical magnetometer, the atoms are spin-polarised using light through the process of optical pumping [26]. In an orientation-based optical magnetometer, the atoms are optically pumped with circularly polarised light. In this case, each atom can be effectively treated as a spin-1/2 particle, even if the ground state of the atom has a total angular momentum $F$ larger than 1/2. The evolution of the atomic spin in a magnetic field is then well described by the Bloch equation for the spin vector $\bm{F}=\left(F_{x},F_{y},F_{z}\right)^{\mathrm{T}}$. Its three components correspond to the expectation values of the angular momentum operators defined along the respective directions, i.e. $F_{\alpha}\coloneqq\mathrm{Tr}\{\hat{F}_{\alpha}\rho\}$ with $\alpha=x,y,z$ for an atomic ensemble being effectively described by a single-atom density matrix $\rho$. On the contrary, in an alignment-based magnetometer [27, 24, 28, 29, 30, 25] the atoms are optically pumped with linearly polarised light. In this case, each atom can be effectively treated as a spin-1 particle [31]. As a result, one has to abandon describing the atomic state with a three-component vector, and instead describe it using rank-2 spherical tensors with five components, which describe how the atomic spin is aligned along certain axes [32].
The purpose of spin noise spectroscopy (SNS) [33] is to characterise the noise properties of a given atomic system and, in particular, the form of the autocorrelation function that noise fluctuations exhibit in the steady-state regime [34]. However, only in the case of orientation-based magnetometers have stochastic noise models been proposed that are capable of explaining the observed noise power spectra when probing unpolarised atomic ensembles [35, 36, 37, 38], also including the effects of spin-exchange collisions [39, 40, 41]. In contrast, such models characterising fully the spin-noise spectra in alignment-based magnetometers are still missing, despite recent promising steps in that direction [42, 43, 44, 45, 46, 47].
@@ II Alignment-based atomic magnetometer
In Fig. 1 we depict the natural spatial configuration of an alignment-based atomic magnetometer [24], which applies to the implementation considered here. A strong magnetic field $\bm{B}_{0}$, directed along $z$, acts perpendicularly to the direction of light propagation, here chosen to be $x$, with light being linearly polarised at an angle $\theta$ to $z$-axis in the $yz$-plane. Both the field and the light interact with an atom and modify its steady state, otherwise induced solely by the relaxation processes [27]. Although we defer the formal description of the atomic steady-state, let us note for now that its form can be conveniently represented by the angular-momentum probability surface [50, 32] (depicted in blue in Fig. 1). As the surface describes the effective polarisation of the steady state, it indicates the rotational symmetries the state possesses. Its shape strongly depends on the polarisation angle $\theta$ of the input light-beam, e.g., resembling a “peanut” for $\theta=0$ or a “doughnut” for $\theta=\tfrac{\pi}{2}$, see Fig. 1(a) and (b), respectively. As such a qualitative description, however, relies solely on symmetry arguments, it applies irrespectively of other properties of the input light, e.g., its detuning.
Importantly, in our experiment the atom is further disturbed by a stochastic field $\bm{B}_{\mathrm{noise}}$, see Fig. 1(c), that induces white noise along the $x$-direction, whose impact can then be monitored by inspecting the polarisation angle $\theta^{\prime}$ of the transmitted beam, while the atomic state is constantly “kicked out” of equilibrium. As we carry out the experiment in conditions in which correlations between distinct atoms may be neglected during the sensing process, the dynamics of the whole ensemble is effectively captured by the evolution of a single atom [26]. Furthermore, due to the effective atomic density being low, the shot noise of the photon-detection process dominates over the atomic projection noise, and the impact of the measurement backaction on the atomic steady-state can be ignored [51, 52]. Moreover, thanks to the presence of a relatively strong magnetic field $\bm{B}_{0}$, given that the experiment is carried out at room temperature, the mechanism of spin-exchange interactions during atomic collisions is irrelevant [53], in contrast to SERF magnetometers [54], that could in principle induce not only classical correlations but also entanglement between individual atoms [55, 56].
@@ II.1 Spherical tensor representation of the atomic steady state
$\rho^{(F)}=\sum_{M,M^{\prime}=-F}^{F}\rho_{MM^{\prime}}^{(F)}\left|F,M\phantom{{}^{\prime}}\right\rangle\!\!\left\langle F,M^{\prime}\right|.$ (1)
However, it is much more convenient to decompose the density matrix in the basis of spherical tensor operators of rank $\kappa=0,1,\dots,2F$, which transform independently in each $\kappa$-subspace under rotations, and hence magnetic fields, in a well-behaved manner. In particular, the atomic state of fixed $F$ in Eq. (1) can be generally written as:
$\rho^{(F)}=\sum_{\kappa=0}^{2F}\sum_{q=-\kappa}^{\kappa}m_{\kappa q}^{(F)}\,{{\mathcal{T}}}^{(\kappa)}_{q}\!(F),$ (2)
where the density matrix is now decomposed into a sum of rank-$\kappa$ components, each constituting a sum (over $q$) of, in principle, non-Hermitian tensor operators ${{\mathcal{T}}}^{(\kappa)}_{q}\!(F)$ (see e.g. [26] and App. A) multiplied by their corresponding complex coefficients $m_{\kappa q}^{(F)}$. For a given fixed $F$, ${{\mathcal{T}}}^{(\kappa)}_{q}$ form a basis and they formally satisfy $\mathrm{Tr}\!\left\{{{\mathcal{T}}}^{(\kappa)}_{q}({{\mathcal{T}}}^{(\kappa^{\prime})}_{q^{\prime}})^{\dagger}\right\}=\delta_{\kappa\kappa^{\prime}}\delta_{qq^{\prime}}$, while it is convenient to also impose ${{\mathcal{T}}}^{(\kappa){\dagger}}_{q}=(-1)^{q}{{\mathcal{T}}}^{(\kappa)}_{-q}$, so that conditions $m_{\kappa q}=(-1)^{q}m^{*}_{\kappa,-q}$ and $\mathrm{Im}\{m_{\kappa 0}\}=0$ ensure then the density matrix (2) is Hermitian. The $m_{00}=1/\sqrt{2F+1}$ coefficient is fixed by the $\mathrm{Tr}\!\left\{\rho\right\}=1$ condition, while the corresponding ($\kappa=0$) tensor operator, $1{{\mathcal{T}}}^{(0)}_{0}=\tfrac{1}{\sqrt{2F+1}}\openone_{2F+1}$, is the only one with a non-zero trace, being invariant under any rotations.
@@
$\rho=\bigoplus_{F}p_{F}\;\rho^{(F)},$ (3)
where $p_{F}$ is the effective fraction of atoms having the total angular momentum $F$. Moreover, one may then focus on the dynamics of only one particular $\rho^{(F)}$ for the $F$-level actually contributing to the light-atom interaction, with $\rho^{(F)}$ being then decomposable just as in Eq. (2). In such a case, as done in what follows, the $(F)$-superscript can be dropped for simplicity.
As the case of a linearly polarised pump is of our interest, see Fig. 1, from symmetry arguments it follows that, independently of the spin-number $F$, only even-$\kappa$ coefficients are modified when interacting with light [57]. Moreover, as we show later, our model predicts no significant coupling between components of different $\kappa$. On the other hand, it is only the orientation ($\kappa=1$) and alignment ($\kappa=2$) components that can be probed by resorting to electric dipole light-atom interactions [26, 57]. Hence, even though the pump in the experiment is relatively strong, so that the atomic state in Eq. (2) reads:
@@
any detection signal obtained by probing the atoms with light is determined by the vector containing the alignment coefficients:
$\textbf{m}=\left(\begin{array}[]{ccccc}m_{2,-2}&m_{2,-1}&m_{2,0}&m_{2,1}&m_{2,2}\end{array}\right)^{\mathrm{T}},$ (5)
whose dynamics must therefore only be tracked.
@@
where $\mathsf{s}_{\theta}\coloneqq\sin\theta$. However, as the strong static field $\bm{B}_{0}$ leads to (Larmor) precession of the atomic state around $z$, see Fig. 1, that is much faster than the timescale of reaching the steady state, i.e. with the Larmor (angular) frequency $\Omega_{L}\coloneqq\gamma_{\mathrm{gmr}}|\bm{B}_{0}|$ much greater than the overall relaxation rate, all the multipoles $m_{2,q\neq 0}$ quickly average to zero, so that according to the secular approximation [26] the steady-state vector (6) simplifies to
@@ II.2.1 Atomic stochastic dynamics
Dissipative non-unitary evolution.
A dissipative evolution of the atomic state (1) is generally described by the Gorini–Kossakowski–Sudarshan–Lindblad equation [58]:
@@
$B_{\mathrm{noise}}(t)=\frac{\Omega_{\mathrm{noise}}}{\gamma_{\mathrm{gmr}}}\,\xi(t)\;\approx\;\frac{\Omega_{\mathrm{noise}}}{\gamma_{\mathrm{gmr}}}\sqrt{\frac{1}{2\pi\Delta f}}\frac{\mathrm{d}W_{t}}{\mathrm{d}t},$ (15)
where $\Omega_{\mathrm{noise}}$ is an effective magnetic noise amplitude given in the units of Larmor frequency, while $\xi(t)$ represents the stochastic process for the noise we generate, see App. C for its further characteristics, which effectively exhibits a constant power spectrum in a frequency range $f\in[\delta\!f,f_{\textrm{cutoff}}]$ (in Hz) with $\delta\!f$ set close to zero to eliminate spurious low-frequency contributions, and some large cut-off $f_{\textrm{cutoff}}$ imposed to prevent the effect of aliasing. Now, as we will deal with processes occurring at (angular) frequencies $\omega\approx\Omega_{L}$, such noise in Eq. (15) can be effectively described as white with the correct rescaling factor $\Delta f\coloneqq f_{\textrm{cutoff}}-\delta\!f\approx f_{\textrm{cutoff}}$—the white noise can be interpreted as the time-derivative, $\frac{\mathrm{d}W}{\mathrm{d}t}$, of the Wiener process, $W_{t}$ [59]. As a result, we can write the (stochastic) time-increment induced by the noise involving the Hamiltonian as
@@
where we define $\omega_{\mathrm{noise}}\coloneqq\Omega_{\mathrm{noise}}^{2}/(2\pi f_{\textrm{cutoff}})$ as the effective noise spectral density, while $\mathrm{d}W_{t}\sim\mathcal{N}(0,\mathrm{d}t)$ is then normally distributed with variance $\mathrm{d}t$, i.e. the Wiener increment [59].
@@ II.2.2 Spherical-tensor representation
Formally, this corresponds to the assumption that in the absence of the induced magnetic-field noise one should differentiate only between dephasing rates along, $\Gamma_{\parallel}\coloneqq\Gamma_{z}$, and perpendicular to, $\Gamma_{\perp}\coloneqq\Gamma_{x}=\Gamma_{y}$, the static field, while keeping the isotropic depolarisation rate, $\Gamma_{\mathrm{iso}}$, as an independent parameter. The effective rates $\bm{\Gamma}$, see also App. A, are then defined by reparametrising the problem as in Eq. (23).
@@
$\mathrm{d}\textbf{m}_{t}$ $=\left(\textbf{A}_{0}+\frac{\textbf{A}_{\mathrm{noise}}^{2}}{2}\right)\textbf{m}_{t}\,\mathrm{d}t+\textbf{A}_{\Phi}(\textbf{m}_{t}-\textbf{m}^{\mathrm{ss}})\,\mathrm{d}t$
@@ II.2.3 Detected signal
We depict the phenomenon of polarisation rotation in Fig. 1(c), i.e. the effect that the angle $\theta$ of the linearly polarised incoming beam is changed to $\theta^{\prime}$ upon leaving the atomic cell. Treating the atomic ensemble as an optically thin sample, the change of the angle, $\Delta\theta\coloneqq\theta^{\prime}-\theta$, obeys then [32, 60]:
@@
where $g_{D}$ is the effective proportionality constant, whereas $\zeta(t)$ denotes the detection noise, which is completely uncorrelated from the magnetic-field noise affecting the atom dynamics in Eq. (24). Moreover, as we assume here the impact of the measurement backaction exerted on the atomic state by the light to be ignorable [51, 52], $\zeta(t)$ shall not exhibit any correlations with any noise exhibited by the atoms.
@@
Figure 2: Alignment-based magnetometer: experimental setup. Linearly polarised light is passed through a polarisation maintaining fibre. Two sets of half-wave plates ($\lambda/2$) and polarising beam splitters (PBS) are used to reduce intensity fluctuations and to change the power of the laser beam, respectively. The laser light propagates along the $x$-direction, while the angle $\theta$ of its linear polarisation is then adjusted by rotating yet another half-wave plate, before it enters the magnetic shield (marked in grey). Within the shield, the light simultaneously pumps and probes the atomic ensemble of caesium atoms contained within a paraffin-coated cell (Cell). The atoms are subject to (strong) static and (relatively weak) noisy magnetic fields: $\bm{B}_{0}$ applied in the $z$-, and $\bm{B}_{\mathrm{noise}}$ applied in the $x$-direction, respectively. Upon interacting with the atoms, the polarisation of light undergoes rotation, which is then measured by passing the output beam through a PBS and performing balanced photodetection (BPD).